Tohban Report 2016-07-27
From RHESSI Wiki
|Start Date:||20 July 2016|
|End Date:||27 July 2016|
|Next Tohban:||Jim McTiernan|
|List all reports|
This week the Sun was particularly active considering that we are entering into quiet zone period of the solar cycle. Starting last Wednesday, AR2567 hosted a total of six C-class flares. The peak of activity occurred on Thursday when this same active region 9 C-class flares and 2 GOES M-class. During the weekend five intense solar limb M-class flares and a total of 15 C-class flares were caused by the same AR. After that "party of solar flares," things began to quiet down this week with just a couple of small flares and a GOES background that fell to a constant B-class.
As a particular situation, the Solar NuSTAR team got solar observation time for yesterday, so during the weekend we prepared RHESSI to support those observations. We expect to hear soon from the NuSTAR team to tell us whether the found something interesting during their observations.
How many GOES flares occurred?
Flares above B, C, M, X class were 26 37 7 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 11 17 3 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 0 0 0 0
There were RHESSI flares/GOES flares 102 / 70 over the time range 19-Jul-16 26-Jul-16
Decimation was normal/vigorous for the full-time period. The SSR is empty. Detectors G3 and G8 were on full-time, including spacecraft night, and detector G9 was turned on starting on Sunday night.
Detectors 3, 8 and 9 are powered on and are recording events. The cold plates, 1 and 2 are at 134.4 and 132.9 K. Detector 9 was turned on 16-207 at 07:06:01 UTC. The cold tip temperatures are 113.1 and 111.1 degrees. Prior to the turn on of D9, the cold tip temperature average was quite stable. The turn on appears to have not affected the temperatures.
We decided to turn detector G9 off in the next days. This is a task for the next Tohban.
Some data gaps listed below.
GAP START TIME GAP END TIME GAP (SEC) 2016-07-20T11:05:00.000 -- 2016-07-20T11:10:00.000 300.00000 GAPS IN APP_ID = 154 (VC1-PMTRAS) WITH PACKET RATE LT 4 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2016-07-20T09:20:00.000 -- 2016-07-20T09:25:00.000 300.00000 NO GAPS IN APP_ID = 102 (VC3-MONITOR RATES) GAP START TIME GAP END TIME GAP (SEC) 2016-07-21T10:00:00.000 -- 2016-07-21T10:05:00.000 300.00000 GAP START TIME GAP END TIME GAP (SEC) 2016-07-25T09:15:00.000 -- 2016-07-25T09:20:00.000 300.00000 2016-07-25T13:45:00.000 -- 2016-07-25T13:50:00.000 300.00000 NO GAPS IN APP_ID = 154 (VC1-PMTRAS) NO GAPS IN APP_ID = 102 (VC3-MONITOR RATES)
Detectors G3, G8 and G9 are powered on and are recording events. No excess large numbers of resets, fast or slow rates at present. All other detectors are off.
|Night time data (fronts)||full-time|
|Night time data (rears)||full-time|
|Require extra passes?||No|
|Requirement for moving pointer?||No|
|Detector problems?||Nominal in the new standard.|